# Fine-grained Defocus Blur Control for Generative Image Models

### v1

@@ 1 Introduction
Our approach is based on training a few-step generative model conditioned on EXIF information (specifically aperture). This model obtains a supervision signal from a differentiable lens model. We train a fast few-step generative model conditioned on EXIF information to generate deep (or shallow) depth-of-field (DoF) images using weak supervision from simply extracting high-quality deep (and shallow) DoF subsets from the dataset. In particular, the shallow DoF images are generated by our novel focus distance transformer, which predicts a focus distance and scale used by a differentiable lens blur model based on Wang et al. [30]. Because focus distances are typically omitted or unreliable in EXIF tags, we show that a weakly supervised approach can be used to learn these directly from shallow DoF images.
By explicitly modeling defocus blur within the lens, our approach enables end-to-end learning of plausible focal distances while manipulating lens properties separately from scene content. We base our generator on improved Distribution Matching Distillation (DMD2) [35] since it can simultaneously learn a fast few-step model to enable user interactivity, and its unpaired DMD and GAN losses are suited for our weak supervision. While these components have been studied in isolation, our key contribution lies in integrating them into a unified, unsupervised training framework that enables precise and intuitive user control over the defocus blur effect — empowering interactive generation of images with controllable DoF.
Our main contribution is a generative framework that enables learning lens properties and allowing precise interactive user control over such lens properties while preserving scene content. Our experiments show clear improvements, enabling such fine-grained user interactivity. Our technical contributions are:
(1) A unified generative pipeline that learns without explicit supervision to model depth of field, which is done with the help of (2) A novel focus distance transformer that learns to predict plausible focus distances and scales, which are used in a physically-inspired thin lens blur model [30]. (3) We show that weak supervision from shallow and deep depth-of-field images can successfully supervise a diffusion model with a lens model, and that this form of supervision can easily be obtained from unlabeled image datasets.
@@ 2 Related Work
Finite apertures produce defocus effects long studied for light-field rendering [2] and adaptive ray-tracing sampling [3, 4, 44], with recent differentiable depth-of-field aware rendering [22]. We draw inspiration from AR-GAN [12], which estimates a depth map and all-in-focus RGB image in a unsupervised way, then integrates over a finite aperture to render defocus. Unlike AR-GAN, we adopt a diffusion framework for text-to-image synthesis and have different learned modules such as the focal-distance model. AR-NeRF [13] uses a similar fully unsupervised approach based on NeRFs [21]. DOF-GS [31] extends 3D Gaussian splatting [15] with a thin-lens circle-of-confusion for 3D reconstruction and defocus. Wang et al. [30] use an implicit neural representation to reconstruct an all-in-focus HDR radiance map and depth map from a time-aperture-focus stack: we use their differentiable thin lens model. Xin et al. [33] use an optimization on a multiplane image to recover an all-in-focus image and defocus map from a dual pixel image. Dr.Bokeh [28] uses a layered scene representation and differentiable occlusion-aware bokeh model to render high-quality bokeh effects. We use their lens model as a plug-and-play component during inference.
@@
A few generative models leverage metadata for conditioning. DiffusionSat [16] generates satellite imagery conditioned on metadata such as GPS, date, cloud cover – some of which, like GPS coordinates, can also be found in EXIF metadata for certain cameras, such as smartphones. EXIF as Language [42] uses contrastive learning to learn connections between image patches and EXIF data for spliced image detection. Camera Settings as Tokens [10], a recent work, trains a LoRA [8] adapter and camera parameter embeddings to condition generation on focal length, aperture, ISO, and exposure. While this approach shows some ability to generate defocus blur based on aperture, it struggles to preserve scene content and fails to consistently adjust blur according to aperture changes. In contrast, our method decouples scene generation from defocus blur, allowing for precise blur control while preserving scene structure. Generative Photography [38], a recent work, uses a text-to-video model to maintain scene consistency across camera settings and learn a camera embedding to capture EXIF data. Their approach, however, relies on collecting paired datasets of images captured at different EXIF tag settings for each metadata type. In contrast, our method is entirely unsupervised, leveraging only unpaired deep and shallow depth-of-field datasets to achieve precise and consistent control over defocus blur while preserving scene content.
@@ 4.1 Evaluation
Figure 4: Qualitative Results. We show our model’s performance across different apertures and prompts. As the aperture increases, background blur decreases while the overall scene remains consistent. Please see the supplemental video for more results.
@@ 5 Results
4-step SDXL (Distilled): We distill the SDXL teacher model into a 4-step generator and train it with DMD2 losses. We train two variants: one that includes a DoF prompt in the text input and another that uses an EXIF projection layer for improved EXIF data processing where we add sinusoidal positional encoding for Aperture and Focal Length and process them using two projection layers and add it to the time embedding. See the supplemental for details regarding the EXIF projection.
@@
Ablations. We ablate key components of our model in Table LABEL:tab:ablations. All ablations are trained on our Shallow and Deep DoF datasets. Rows 1 and 2 exclude the lens model and instead use aperture and focal length as additional text inputs to control blur behavior. These configurations show reduced Blur Consistency, indicating that representing aperture solely as text does not provide sufficient information to consistently simulate blur changes. Notably, Row 2 outperforms Row 1, demonstrating the benefit of pretraining on the Deep DoF dataset. This pretraining provides our generator $\mathbf{G}$ with a strong prior for generating all-in-focus images during fine-tuning.
In row 3, the addition of the lens model improves Content Consistency when changing aperture and enhances the model’s ability to apply blur consistently at a given aperture. Row 3 removes the focal distance model and replaces it with mean depth, which compared to the full model gives worse visual quality (FID) and worse blur monotonicity as the focus distance scale is not learned, causing unintended out-of-focus blur and limiting the lens model’s ability to increase blur beyond a certain point. Finally, in row 4, adding the focus distance model allows the lens model to better focus on salient regions, reducing out-of-focus artifacts.
@@ 6 Limitations, Future Work, Conclusion
Appendix A Comparisons to our model
@@
Appendix B Controllability of defocus blur in image generation
@@
Appendix C Deep and Shallow Depth-of-Field Datasets
Figure A4: Deep and Shallow DoF Datasets. Images shown are selected using our dataset filtering approach mentioned in Sec. C. After filtering, the Deep DoF dataset primarily consists of all-in-focus images, while the Shallow DoF dataset includes images with defocus blur, emphasizing a specific object of interest. We use these datasets to train our model.
@@ Appendix C Deep and Shallow Depth-of-Field Datasets
Appendix D Human Study
@@
Appendix E Implementation Details
@@ E.2 Evaluation Metrics
E.3 4-step SDXL (distilled) with EXIF
@@
Figure A6: Human Studies. Preference rate for selecting our method over the baselines (Section  D).

### v2

@@ 1 Introduction
Our approach is based on training a few-step generative model conditioned on EXIF information (specifically aperture). This model obtains a supervision signal from a differentiable lens model. We train a fast few-step generative model conditioned on EXIF information to generate deep (or shallow) depth-of-field (DoF) images using weak supervision from simply extracting high-quality deep (and shallow) DoF subsets from the dataset. In particular, the shallow DoF images are generated by our novel focus distance transformer, which predicts a focus distance and scale used by a differentiable lens blur model based on Wang et al. [32]. Because focus distances are typically omitted or unreliable in EXIF tags, we show that a weakly supervised approach can be used to learn these directly from shallow DoF images. By explicitly modeling defocus blur within the lens, our approach enables end-to-end learning of plausible focal distances while manipulating lens properties separately from scene content. We base our generator on improved Distribution Matching Distillation (DMD2) [38] since it can simultaneously learn a fast few-step model to enable user interactivity, and its unpaired DMD and GAN losses are suited for our weak supervision. While these components have been studied in isolation, our key contribution lies in integrating them into a unified, unsupervised training framework that enables precise and intuitive user control over the defocus blur effect — empowering interactive generation of images with controllable DoF.
Our main contribution is a generative framework that enables learning lens properties and allowing precise interactive user control over such lens properties while preserving scene content. Our experiments show clear improvements, enabling such fine-grained user interactivity. Our technical contributions are: (1) A unified generative pipeline that learns without explicit supervision to model depth of field, which is done with the help of (2) A novel focus distance transformer that learns to predict plausible focus distances and scales, which are used in a physically-inspired thin lens blur model [32]. (3) We show that weak supervision from shallow and deep depth-of-field images can successfully supervise a diffusion model with a lens model, and that this form of supervision can easily be obtained from unlabeled image datasets.
@@ 2 Related Work
Finite apertures produce defocus effects long studied for light-field rendering [2] and adaptive ray-tracing sampling [4, 5, 47], with recent differentiable depth-of-field-aware rendering [24]. AR-GAN [14] estimates a depth map and all-in-focus RGB image in an unsupervised manner and integrates over a finite aperture to render defocus; unlike it, we adopt a diffusion framework for text-to-image synthesis with learned modules such as the focal-distance model. AR-NeRF [15] uses a similar unsupervised approach based on NeRFs [23], while DOF-GS [33] extends 3D Gaussian splatting [17] with a thin-lens circle-of-confusion for 3D reconstruction and defocus. Wang et al. [32] reconstruct an all-in-focus HDR radiance map and depth map from time-aperture-focus stacks, whose differentiable thin lens model we use, while Xin et al. [36] recover all-in-focus images and defocus maps from dual-pixel images via multiplane optimization. Dr.Bokeh [30] uses a layered scene representation with a differentiable occlusion-aware bokeh model, which we adopt as a plug-and-play lens component during inference. DC2 [3] enables post-capture defocus control by fusing dual-camera inputs with different fixed apertures but requires multi-camera hardware. DiffCamera [34] performs arbitrary post-capture refocusing using a diffusion transformer trained on simulated multi-focus data, whereas our method learns defocus from in-the-wild data and provides fine-grained control over focus and aperture.
@@
Several models leverage metadata for conditioning. DiffusionSat [18] generates satellite imagery conditioned on metadata such as GPS, date, cloud cover, some of which also appear in EXIF metadata for cameras. EXIF as Language [45] uses contrastive learning to connect image patches with EXIF data for spliced image detection. Camera Settings as Tokens [12] trains a LoRA [10] adapter and camera parameter embeddings to condition generation on focal length, aperture, ISO, and exposure. but struggles to preserve scene content and consistently control blur. In contrast, our method decouples scene generation from defocus blur for precise control. Generative Photography [41], learns camera embeddings using paired videos with varying EXIF settings, whereas our method is entirely unsupervised, leveraging only unpaired deep and shallow depth-of-field data. A concurrent work, Bokeh Diffusion [8], conditions a diffusion model on a defocus parameter for scene-consistent bokeh control but relies on synthetic blur for supervision.
@@ 4.1 Evaluation
Figure 4: Qualitative Results. We show our model’s performance across different apertures and prompts. As the aperture increases, background blur decreases while the overall scene remains consistent. Please see the website videos for more results.
@@ 5 Results
4-step SDXL (Distilled): We distill the SDXL teacher model into a 4-step generator and train it with DMD2 losses. We train two variants: one that includes a DoF prompt in the text input and another that uses an EXIF projection layer for improved EXIF data processing where we add sinusoidal positional encoding for Aperture and Focal Length and process them using two projection layers and add it to the time embedding. See the supplemental section F.3 for details regarding the EXIF projection.
@@
Ablations. We ablate key components of our model in Table LABEL:tab:ablations. All ablations are trained on our Shallow and Deep DoF datasets (except row 5). Rows 1 and 2 exclude the lens model and instead use aperture and focal length as additional text inputs to control blur behavior. These configurations show reduced Blur Consistency, indicating that representing aperture solely as text does not provide sufficient information to consistently simulate blur changes. Notably, row 2 outperforms row 1, demonstrating the benefit of pretraining on the Deep DoF dataset. This pretraining provides our generator $\mathbf{G}$ with a strong prior for generating all-in-focus images during fine-tuning.
In row 3, the addition of the lens model improves Content Consistency when changing aperture and enhances the model’s ability to apply blur consistently at a given aperture. Row 3 removes the focal distance model and replaces it with mean depth, which compared to the full model gives worse visual quality (FID) and worse blur monotonicity as the focus distance scale is not learned, causing unintended out-of-focus blur and limiting the lens model’s ability to increase blur beyond a certain point. Finally, in row 4, adding the focus distance model allows the lens model to better focus on salient regions, reducing out-of-focus artifacts. Row 5 shows pretraining the generator on Deep-DoF and freezing it, followed by finetuning only on Shallow DoF, leads to worse all-in-focus image generation.
@@ 6 Limitations, Future Work, Conclusion
Appendix A Video
We provide videos on our webpage (link) showing the controllability of defocus blur using our model. We also include qualitative examples demonstrating defocus control in generated images for several prompts.
Appendix B Comparisons to our model
@@
Appendix C Controllability of defocus blur in image generation
@@
Appendix D Deep and Shallow Depth-of-Field Datasets
Figure A4: Deep and Shallow DoF Datasets. Images shown are selected using our dataset filtering approach mentioned in Sec. D. After filtering, the Deep DoF dataset primarily consists of all-in-focus images, while the Shallow DoF dataset includes images with defocus blur, emphasizing a specific object of interest. We use these datasets to train our model.
@@ Appendix D Deep and Shallow Depth-of-Field Datasets
Appendix E Human Study
@@
Appendix F Implementation Details
@@ F.2 Evaluation Metrics
F.3 4-step SDXL (distilled) with EXIF
@@
Figure A6: Human Studies. Preference rate for selecting our method over the baselines (Section  E).